# Jacobian rings and the infinitesimal Torelli Theorem

### v1

@@ 1 Introduction
The $g_{\Gamma}(f)$’s in fact are useful generators of the first component $J_{\Delta,f}^{1}$ of the jacobian ideal. These formulas heavily rely on the nondegeneracy of $f$ which could be translated in an algebraic independence between certain Laurent polynomials $F_{0},...,F_{n}$ (see theorem 2.4). A deficit is that the above generators of $U_{f,k}$ are in general not linearly independent for $k\geq 3$.
@@
with $\{F^{k}\}$ the Hodge filtration. This map is known to be holomorphic. The differential $d\mathcal{P}_{f}^{k}$ factors as follows
@@
Our idea is strainghtforward: Compute the kernel $\ker(d\mathcal{P}_{B,f}^{k})$ and compare it with $\ker(\kappa_{f})$, which we explicitly computed in the preceeding article [Gie22], to finally deduce on $\ker(\Phi_{f|{Im(\kappa_{f})}}^{k})$. The key of our results is the following theorem, which allows a computational approach (theorem 8.4)
@@
Let $\Delta$ be an $n$-dimensional lattice polytope with $\operatorname{Int}(\Delta)\cap M\nsubseteq\textrm{ hyperplane}$. Then
@@
Given $n=3$ assume $\operatorname{Int}(\Delta)\cap M\nsubseteq\textrm{plane}$. Then
@@
Concerning the cokernel $\operatorname{coker}(\kappa_{f})$ the situation might get more complicated. In ([Mav03]) there were made some efforts for $Y_{f}$ a K3 hypersurface (so called non-polynomial deformations), realizing these deformations as deformations of complete intersections in a single big toric variety, but is seems unlikely that this representation is useful in order to compute the missing vector space $\ker(\Phi_{f|{\operatorname{coker}(\kappa_{f})}})$. In ([BrGr12]) there were made some efforts to combine the Hodge components and their computation to the generic Picard number of the hypersurface.
@@ 4 Hodge and mixed Hodge structures
([Del75], see also [Dan78, §14])
For $Y$ an $n$-dimensional quasismooth hypersurface in a simplicial projective toric variety $\mathbb{P}$ the cohomology groups $H^{k}(Y,\mathbb{Q})$ are equipped with a Hodge structure of weight $k$ and we have
@@
For $\Delta$ an $n$-dimensional lattice polytope, $f\in U_{reg}(\Delta)$ and $Z_{f}=\{f=0\}\subset T$ the cohomology groups
$H^{k}(Z_{f},\mathbb{Q})$
carry a mixed Hodge structure of weight $k$.
@@ 8 Explicit description of $d\phi_{f}$
Let $\Delta$ be an $n$-dimensional lattice polytope with $\operatorname{Int}(\Delta)\cap M\nsubseteq\textrm{ hyperplane}$. Then
@@
but $h_{\Gamma}\cdot g_{\Gamma}(f)\notin L(\Delta)$ for several $\Gamma$’s. We start with two Lemmas: Suppose given an element of the kernel as in (8.8). Let
@@
For simplicity assume in the following $k=1$. Choose one aribitrary distinguished facet $\Gamma_{1}$ and restrict to those $(m,v)\in P_{\Gamma}$ with $\langle m+v,n_{1}\rangle<\operatorname{Min}_{\Delta}(n_{1})$ and without restriction $\langle m+v,n_{1}\rangle=\textrm{ minimal}$ for some fixed facet $\Gamma_{1}$ with inner facet normal $n_{1}$. We adopt the obvious notation $P_{i}$, $n_{i}$ and $h_{i}$.
@@
by (8.4), a contradiction.
To (benumi): Here
$\langle\operatorname{Int}(\Delta)\cap M,n_{s}\rangle\geq 0,\quad\quad\langle\operatorname{Int}(\Delta)\cap M,n_{1}\rangle\geq 0,$ (27)
since else
$\big(\operatorname{Int}(\Delta)\cap M\big)+m_{i}\notin\Delta\cap M,\quad\textrm{for }i=1,s.$
If the equation for $m_{s}$ is an equality, then apparently $\operatorname{Int}(\Delta)\cap M$ is contained in an hyperplane, a contradiction. Take $\tilde{m}\in\operatorname{Int}(\Delta)\cap M$ with $\langle\tilde{m},n_{s}\rangle>0$. If $\langle\tilde{m},n_{1}\rangle=0$ then
$\tilde{m}+m_{s}\in\operatorname{Int}(\Delta)\cap M$ (28)
and
$\langle m+\tilde{m}+m_{s},n_{1}\rangle=\langle m+m_{s},n_{1}\rangle<\operatorname{Min}_{\Delta}(n_{1})$
but $m+\tilde{m}+m_{s}\in R_{\Gamma_{1}}$, a contradiction. If $\langle\tilde{m},n_{1}\rangle>0$ then
$\langle\tilde{m}+k\cdot m,n_{1}\rangle=0\textrm{ for some }k\geq 1$
Replace $\tilde{m}$ by $\tilde{m}+k\cdot m$ and continue as above.
To (benumi): Take $k:=k(m)\in\mathbb{N}_{\geq 1}$. There is $(m_{s},v_{s})\in P_{s}$ with

### v2

@@ 1 Introduction
The $g_{\Gamma}(f)$’s in fact are useful generators of the first component $J_{\Delta,f}^{1}$ of the jacobian ideal. These formulas rely on the nondegeneracy of $f$ which could be translated in an algebraic independence between certain Laurent polynomials $F_{0},...,F_{n}$ (see theorem 2.4). A deficit is that the above generators of $U_{f,k}$ are almost never linearly independent for $k\geq 3$.
@@
with $\{F^{k}\}$ the Hodge filtration. This map is known to be holomorphic. The differential $d\mathcal{P}_{B,f}^{k}$ factors as follows
@@
Our idea is strainghtforward: Compute the kernel $\ker(d\mathcal{P}_{B,f}^{k})$ and compare it with $\ker(\kappa_{f})$, which we explicitly computed in the preceeding article [Gie22], to finally deduce on $\ker(\Phi_{f|{Im(\kappa_{f})}}^{k})$. The key of our results is the following theorem (theorem 8.4)
@@
Let $\Delta$ be an $n$-dimensional lattice polytope such that $\operatorname{Int}(\Delta)\cap M$ is not contained in an hyperplane. Then
@@
Given the assumptions of theorem 1.2
@@
Concerning the cokernel $\operatorname{coker}(\kappa_{f})$ the situation might get more complicated, see also [Gie22, Cor.7.1]. In ([Mav03]) there were made some efforts for $Y_{f}$ a K3 hypersurface (so called non-polynomial deformations), realizing these deformations as deformations of complete intersections in a single big toric variety, but is seems unlikely that this representation is useful in order to compute the missing vector space $\ker(\Phi_{f|{\operatorname{coker}(\kappa_{f})}})$. Besides in ([BrGr12]) there were made some efforts to combine the Hodge components and their computation to the generic Picard number of the hypersurface.
@@ 4 Hodge and mixed Hodge structures
([Del75])
For $\Delta$ an $n$-dimensional lattice polytope, $f\in U_{reg}(\Delta)$, the cohomology groups $H^{k}(Y,\mathbb{Q})$ are equipped with a Hodge structure of weight $k$ and we have
@@
Besides $H^{k}(Z_{f},\mathbb{Q})$ carry a mixed Hodge structure of weight $k$.
@@ 8 Explicit description of $\ker(d\phi_{f})$
Let $\Delta$ be an $n$-dimensional lattice polytope such that $\operatorname{Int}(\Delta)\cap M$ is not contained in an hyperplane. Then
@@
$w+v\in\big(\operatorname{Int}((k+1)\cdot\Delta)\cup\operatorname{Int}((k+1)\cdot\Gamma)\big)\cap M,$ (21)
@@
but $h_{\Gamma}\cdot g_{\Gamma}(f)\notin L(\Delta)$ for several $\Gamma$’s.
Remark 8.5.
Given $\Delta$ with $(0,..,0)\in\operatorname{Int}(\Delta)\cap M$ and $g_{\Gamma}(f)\cdot x^{w}\in\ker(d\phi_{f}^{k})$ assume that we replace $M$ by some $p\in M$ such that the new zero point p-p is again an interior lattice point of $\Delta$. Then $g_{\Gamma}(f)\cdot x^{w}$ gets well transformed to $g_{\tilde{\Gamma}}(f)\cdot x^{w}$, where $g_{\tilde{\Gamma}}(f)\cdot x^{-p}=g_{\Gamma}(f)$, with equivalent relation 8.4. This is basically because $p\in\operatorname{Int}(\Delta)\cap M$ and due to the relation 8.4 again.
We start with two Lemmas: Suppose given an element of the kernel as in (8.8). Let
@@
For simplicity assume in the following $k=1$. Choose one aribitrary distinguished facet $\Gamma_{1}$ and restrict to those $(m,v)\in P_{\Gamma}$ with $\langle m+v,n_{1}\rangle<\operatorname{Min}_{\Delta}(n_{1})$ and without restriction $\langle m+v,n_{1}\rangle=\textrm{ minimal}$ for some fixed facet $\Gamma_{1}$. We adopt the obvious notation $P_{i}$, $n_{i}$ and $h_{i}$.
@@
a contradiction.
To (benumi): Obviously we have $n_{1}\neq-n_{s}$ since $m_{s}\in\Gamma_{s}$ and $\langle m_{s},n_{1}\rangle<0$. Take $F:=H_{n_{1},0}\cap H_{n_{s},0}$ and the linear subspace $E:=\langle F,m\rangle$. Since $n_{1}\neq-n_{s}$ the subspace $F$ is (n2)(n-2)-dimensional and $E$ contained in an hyperplane. With the hypotheses of the theorem take $\tilde{m}\in\operatorname{Int}(\Delta)\cap M$ with $\tilde{m}\notin E$. Then $\langle\tilde{m},n_{1}\rangle\geq 0$ since $m\in\Gamma_{1}$ and if $\langle\tilde{m},n_{s}\rangle=0$ then $\tilde{m}+k\cdot m\in F$, that is $\tilde{m}\in E$, a contradiction. If $\langle\tilde{m},n_{s}\rangle<0$, then $m_{s}\in\Gamma_{s}$ yields a contradiction. Thus $\langle\tilde{m},m_{s}\rangle>0$. Choose $k\in\mathbb{N}$ with $\langle\tilde{m}+k\cdot m,n_{1}\rangle=0$. Then again $\langle\tilde{m}+k\cdot m,n_{s}\rangle>0$, $m_{s}+(\tilde{m}+k\cdot m)\in\operatorname{Int}(\Delta)\cap M$ and
$m_{s}+(\tilde{m}+k\cdot m)+m\in R_{1}$
but $\langle m_{s}+(\tilde{m}+k\cdot m)+m,n_{1}\rangle=\langle m_{s}+m,n_{1}\rangle<\operatorname{Min}_{\Delta}(n_{1})$, a contradiction.
To (b) : Take $k:=k(m)\in\mathbb{N}_{\geq 1}$. There is $(m_{s},v_{s})\in P_{s}$ with